# The Promise of Time-Series Foundation Models for Agricultural Forecasting: Evidence from Marketing Year Average Prices

### v1

@@ 5.5 Limitations and Future Research
Data scope. Our sample is limited to four commodities with relatively liquid futures markets and established USDA forecasting benchmarks. Extending to livestock and dairy prices would test performance on commodities with different biological production cycles (gestation periods vs. growing seasons) and storage constraints (perishability vs storability). Specialty crops (fruits, vegetables, nuts) present additional challenges: limited or no futures markets, higher price volatility from weather sensitivity, and stronger regional variation in production and pricing. International commodity prices from FAO or World Bank databases would assess whether foundation models generalize across different policy regimes, trade patterns, and market structures. Higher-frequency data (daily futures prices) would test short-horizon forecasting capabilities relevant for trading and risk management.
@@ 6 Conclusion
This paper provides the first systematic evaluation of time series foundation models for agricultural price forecasting. We compare five foundation models (Chronos, Chronos-2, TimesFM 2.5, Time-MoE, Moirai-2) against 12 baselines spanning traditional time series, machine learning, and deep learning approaches, using USDA ERS data for corn, soybeans, wheat, and cotton from 1997-2025.
Our central finding is that zero-shot foundation models substantially outperform all alternatives without requiring domain-specific training, presenting a potential paradigm shift in this area. Time-MoE achieves 45.4% improvement over USDA’s operational forecasts, with particularly large gains on corn (52.9%) and soybeans (55.2%). All five foundation models rank in the top five positions, while deep learning models trained from scratch rank 10th-17th due to severe overfitting with limited training data. Smaller mixture-of-experts architectures (50M parameters) outperform larger dense transformers (200M parameters), demonstrating that architectural efficiency and massive pre-training scale matter more than raw parameter count.
These findings have direct policy implications for Farm Bill program administration, where improved MYA forecasts could affect billions of dollars in payments to U.S. farmers. More broadly, our results demonstrate that foundation models offer a viable path forward for economic forecasting in data-scarce domains where traditional machine learning has historically struggled. The strong performance despite using only historical prices – outperforming USDA’s futures-based methodology – suggests that historical basis adjustments may introduce systematic bias, pointing to potential improvements in operational forecasting practices.

### v2

@@ 5.5 Limitations and Future Research
Data scope. Our sample focuses on four major field crops with liquid futures markets and established USDA forecasting benchmarks. The broader USDA forecasting system encompasses livestock production (beef, pork, poultry), dairy prices, and specialty crops, each presenting distinct forecasting challenges (Isengildina-Massa et al., 2024). Extending our evaluation to livestock and dairy would test foundation model performance on commodities with different biological production cycles (gestation periods vs. growing seasons) and storage constraints (perishability vs. storability). Specialty crops (fruits, vegetables, nuts) present additional challenges: limited or no futures markets, higher price volatility from weather sensitivity, and stronger regional variation in production and pricing. International commodity prices from FAO or World Bank databases would assess whether foundation models generalize across different policy regimes, trade patterns, and market structures. Higher-frequency data (daily futures prices) would test short-horizon forecasting capabilities relevant for trading and risk management.
@@ 6 Conclusion
This paper provides the first systematic evaluation of time series foundation models for agricultural price forecasting, comparing five foundation models against 12 baselines using USDA ERS data for corn, soybeans, wheat, and cotton from 1997-2025.
For monthly price forecasting, foundation models dominate performance rankings, occupying all top five positions. Time-MoE achieves the lowest mean absolute error, followed by Chronos, Chronos-2, TimesFM 2.5, and Moirai-2. Deep learning models trained from scratch rank 10th-17th despite extensive hyperparameter optimization, while the Naive model (rank 6) outperforms all deep learning approaches. For MYA forecasting against USDA ERS benchmarks, foundation models achieve strong performance on major commodities despite using only historical prices while USDA incorporates forward-looking futures market information. On recent data, Time-MoE delivers substantial improvements on wheat and corn, while TimesFM achieves modest gains on soybeans. These findings challenge the long-standing belief that simple models forecast best in agricultural markets.
These findings have direct policy implications for Farm Bill program administration, where improved SAP forecasts could affect billions of dollars in payments to U.S. farmers. Beyond farm policy applications, our findings have implications for commodity market participants. Prior research notes that futures-based forecasts provide viable alternatives to USDA forecasts for private sector users. Our results suggest that foundation models, requiring only historical prices and no specialized market data, offer an accessible forecasting tool for agricultural businesses, commodity traders, and portfolio managers seeking reliable price projections.
More broadly, our findings have implications beyond agricultural forecasting. The data characteristics that challenge agricultural price prediction (limited samples, structural breaks, strong seasonality, high volatility) are common across economic and business forecasting domains. Foundation models’ ability to outperform specialized benchmarks despite these challenges suggests their potential applicability to other data-scarce forecasting problems, including financial time series, macroeconomic indicators, and business demand forecasting. The synthetic benchmark results (Appendix D) further demonstrate that these advantages stem from genuine pattern learning rather than domain-specific memorization, as foundation models maintain top performance on synthetically generated data with zero contamination risk. Our results demonstrate that foundation models offer a viable path forward for economic forecasting in data-scarce domains where traditional machine learning has historically struggled.